# Guided filtering and smoothing for infinite-dimensional diffusions

### v1

@@ 3.1. Tempering and moving steps
Even though the proposal kernel introduced in Equation (32) defines a suitable, likelihood-informed importance sampling distribution, the guided particle filter might still face weight degeneracy due to the high dimensionality of the state space. Since this is a common issue in particle-based inference, many solutions to overcome it have been proposed and studied in the literature. For the purpose of our numerical experiments in Section 5, we introduce two such solutions - tempering and moving (or particle rejuvenation) steps - into the guided particle filter. This follows the suggestions made in [Llopis et al., 2018], where the filtering problem was considered in a similar setup to ours. The following brief exposition is based on that work.
@@
for a set of inverse temperatures $10=\psi_{0}<\ldots<\psi_{L}=$ 1. Notably, we have absolute continuity of $\Pi^{l+1}$ with respect to $\Pi^{l}$ with R.N. derivative proportional to $\Lambda(X)^{\psi_{l+1}-\psi_{l}}$. Instead of targeting $\Pi$ directly by $\mathbb{Q}$, one progressively steps through this sequence of artificial distributions, starting from $\mathbb{Q}$ to target the flattened $\Pi^{1}$ and ending up at targeting $\Pi$ with proposals drawn from $\Pi^{L-1}$.
@@
After $\Pi^{l+1}$ has been determined, particles are resampled based on the weights $\Lambda(X)^{\psi_{l+1}-\psi_{l}}$. Subsequently, to increase particle diversity, each particle is moved by a few steps of an MCMC scheme that targets $\Pi^{l+1}$. This can be done by proposing independent samples from $\mathbb{Q}$ and accepting/rejecting the proposals in a Metropolis-Hastings step based on the likelihood in Equation (36). Alternatively, proposals can be localised based on the preconditioned Crank-Nicolson scheme, see Section 4 for details. The complete tempering and move procedure is summarised in Algorithm 2. It is to be carried out at each observation time $t_{i}$ in place of Steps $2$ and $3$ in Algorithm 1.
Input: Samples $X^{(j)}$ of $\mathbb{Q}=\Pi_{i-1}\otimes\mathcal{L}^{g_{i}}$ with weights $\Lambda(X^{(j)})$, hyperparameters $\alpha$ and $N$
@@ 3. Move particles:
For ease of exposition, the MCMC move step in Algorithm 2 is presented using independent proposals drawn from $X^{\prime}\sim\mathbb{Q}$. These proposals can be localised in Wiener space by using the preconditioned Crank-Nicolson scheme as is further explained in Section 4. For practical purposes, this implies having to store at the observation time $t_{i}$ for each particle $X^{(j)}$ the sample of the Wiener process $W$ that corresponds to simulating the sample path $X^{(j)}_{[t_{i-1},t_{i}]}$ following the dynamics in (33). Note that this is still carried out in an online setting since the stored Wiener paths relate to the evolution on $[t_{i-1},t_{i}]$ only and may hence be discarded after each complete filtering step.
@@ 4. Guided smoothing via Metropolis-Hastings sampling
To localise the proposals drawn from $\mathcal{L}^{g}(X)$ we introduce random-walk-type proposals in the Wiener space by adapting a preconditioned Crank-Nicolson (pCN) scheme (see [Neal, 1999], [Beskos et al., 2008], [Cotter et al., 2013]) as follows.
@@
Input: Parameters $A,~{}F,~{}Q$ and $x_{0}$, observations $y$, iterations $N,N_{0}$, step size $\beta$
@@ 2. Save current state $X$
Since $W$ is a cylindrical Wiener process on $H$, it cannot be treated directly as an $H$-valued process. However, $W$ can be defined as a Wiener process taking values in a larger Hilbert space $H^{\prime}$ such that $H\hookrightarrow H^{\prime}$ is embedded in a Hilbert-Schmidt way. Note that such an embedding always exists, see [Hairer, 2009], Section 4.4 for details. The solution map $\Gamma$ in (37) is then to be understood as a measurable mapping $\Gamma:H\times C([0,T];H^{\prime})\to C([0,T];H)$.
@@
Input: Parameters $A,~{}F,~{}Q$ and $\rho$, observations $y$, iterations $N,N_{0}$, step sizes $\beta,\beta_{0}$
@@ 4. Save current state $(V,X_{0},\theta)$
Remark 4.4.
If $A$ and $Q$ are independent of the unknown parameter $\theta$, one only needs to solve the backwards ODEs $(U_{t})_{t\in[0,T]}$ and $(V_{t})_{t\in[0,T]}$ in (29) and (30) once. This severely eases the computational burden of Algorithm 5 for the highly relevant case that the parameter $\theta$ of interest parametrises only the non-linearity $F=F_{\theta}$.
@@ 5.1. Stochastic Amari equation
with non-linearity $F:L^{2}(D)\to L^{2}(D)$ defined by
@@
with respect to the Fourier basis functions $(e_{l})_{l=1}^{\infty}$ on $L^{2}(D)$. The operator $Q$ is then the Hilbert-Schmidt integral operator defined by the Matérn kernel on $D$, see for example [Borovitskiy et al., 2020], Theorem 5.
@@ 5.2. Filtering
Experiments are carried out on data generated by forward simulating the solution $X$ to Equation (43) on the spatio-temporal domain $[0,T]\times D=[0,20]\times[-10\pi,10\pi]$ with step sizes $\Delta_{t}=0.02$ and $\Delta_{x}=20\pi/2^{8}$.
@@ Methods
    The guided particle filter as proposed in [Llopis et al., 2018]. This method differs to ours in the proposal distribution $\mathbb{Q}$ of Eq. (32). The proposals chosen in [Llopis et al., 2018] are defined by solving the guided process (33) with guiding term
@@
    The unscented Kalman filter (UKF) introduced in [Wan and Van Der Merwe, 2000]. The UKF has been proposed for the filtering problem of an Amari model in [Schiff and Sauer, 2007], where it has been shown to perform well in the setting of a discretisation of Equation (43).
@@
Hence, once equilibrium is reached, the non-linear part of the dynamical system vanishes, rendering a temporal discretisation more forgiving. As can be observed in Figure 3(d), with increasing time lag between observations, the performance of the UKF worsens drastically to the point of diverging in the case of $n=5$.
A comparison of the two particle filter methods in Figure 3(a) shows that GPF-II with proposals based on the guiding function in (49) slightly outperforms our proposed GPF-I in the dense observation setting. This may be explained yet again by the models convergence to an equilibrium state. Since the guiding function defined in (49) is derived based on the choice $\mathop{}\!\mathrm{d}Z_{t}=\sqrt{Q}\mathop{}\!\mathrm{d}W_{t}$ for the auxiliary process $Z$, once equilibrium is reached and $X$ behaves following (50), such proposals are almost locally optimal. Hence, GPF-II starts outperforming GPF-I around the observation times just before $t=8.0$ when an equilibrium state is ‘fully’ reached the first time. On the other hand, when observations are sparse, the quality of the proposals for the GPF-I at states away from equilibrium lead to a better performance in the cases of $n=10$ and $n=5$. Interestingly, as can be seen in Figure 3(b), the performance of GPF-I in this experiment increases as observations become more sparse. This coincides with findings in [Llopis et al., 2018] for GPF-II in the context of a different dynamical system. In our experiment, this increase in performance is drastic enough for GPF-I in the $n=5$ case to be the best performing method over all observation schemes.
@@
For the second filtering experiment, the comparison between relative filtering errors of the three methods can be found in Figure 5. In this parametrisation with $\delta=0.5$, the system does not converge to an equilibrium; instead, it exhibits traveling wave behaviour. Expectedly, the proposed GPF-I outperforms GPF-II over all three observation schemes. We also see the UKF performing worse than the particle filter methods, even in the dense observation scheme. Moreover, the large time steps between observations plus the non-linear dynamics of the traveling waves states result in significantly more divergent behaviour in the case of $n=5$ observations. On the other hand, GPF-II performs only slightly worse as observation numbers decrease, whereas GPF-I remains robust to the sparsity of observations.
@@ 5.3. Smoothing and parameter estimation
[…]
@@
The parameters $\{\eta,\zeta,A,\delta\}$ are updated individually in a Gibbs sampling manner. Proposals for each parameter are chosen via an adaptive scaling Metropolis scheme (see e.g. [Andrieu and Thoms, 2008]) as follows. At iteration $j$, given a step size $S^{\eta}_{j}$, say we propose $\eta_{j+1}$ via
@@ Results
[…]
@@
[…]
@@
Let first $t\in[0,t_{1}]$ and denote by $K$ the infinitesimal space-time generator of the Markov process $X$. Following (51), it holds that $Kh=0$. From this and the Fréchet differentiability of $h$ in $x$ it follows with Lemma 3.3 in [Pieper-Sethmacher et al., 2025a] that
@@
This, jointly with $E^{h}_{t}=h(t,X_{t})/C^{h},~{}t\in[0,t_{1}],$ shows (53) on $[0,t_{1}]$. Now, let $t\in(t_{i-1},t_{i}],i\geq 2,$ and assume (53) holds for all $[0,t_{i-1}].$
Again, by the definition of $E^{h}$, Equation (51) and Lemma 3.3 in [Pieper-Sethmacher et al., 2025a] we have that
@@
Step Three. The general claim follows by repeating the backwards recursion of the first two steps to obtain $g(t,x)$ on the complete interval $[0,t_{n}].$
@@
Hence, denoting by $K$ the infinitesimal space-time generator of $X$, it follows from [Manca, 2009], Theorem 4.1 that $Kg(t,x)=\langle F(t,x),\mathop{}\!\mathrm{D}_{x}g(t,x)\rangle$. Lemma 3.3 of [Pieper-Sethmacher et al., 2025a] therefore gives that
@@
Consequently, plugging in $E^{g}$ as in (23) and applying the integration by parts for semimartingales we have on $[0,t_{1})$ that
@@
This shows that $\bar{E}^{g}_{t}:=E^{g}_{t}/E^{g}_{0}=\mathcal{E}(M^{g})_{t}$ and hence (55) for all $t\in[0,t_{1}]$. The extension onto the complete time interval $[0,T]$ follows by the same inductive argument as in the proof of Theorem 2.4. Lastly, to show that $\mathcal{E}(M^{g})$ is a proper martingale, it suffices to note that $x\mapsto G(t,x)$ is Lipschitz continuous in $x$, uniformly on $[0,T]$. The martingale property of $\mathcal{E}(M^{g})$ then follows from [Pieper-Sethmacher et al., 2025a], Lemma C.1. ∎
@@
This shows (57) and uniqueness of the mild solution follows from uniqueness of the infinite-dimensional forward Ricatti equation, see for example [Burns and Rautenberg, 2015], Theorem 3.6 as well as [Bensoussan et al., 2007], Chapter 2.2 and references within.
@@ B.1. Proofs of Section 5
Given $Ax=-x$, it holds that $(S_{t})_{t}$ is the semigroup given by $S_{\Delta_{i}}=\exp(-\Delta_{i})$. Hence, following the definition of $Q_{t}$ given in (16), we have

### v2

@@ 3.1. Tempering and moving steps
Even though the proposal kernel introduced in Equation (31) defines a suitable, likelihood-informed importance sampling distribution, the guided particle filter might still face weight degeneracy due to the high dimensionality of the state space. Since this is a common issue in particle-based inference, many solutions to overcome it have been proposed and studied in the literature. For the purpose of our numerical experiments in Section 5, we introduce two such solutions - tempering and moving (or particle rejuvenation) steps - into the guided particle filter. This follows the suggestions made in [44], where the filtering problem was considered in a similar setup to ours. The following brief exposition is based on that work.
@@
for a set of inverse temperatures $0=\psi_{0}<\ldots<\psi_{L}=1$. Notably, we have absolute continuity of $\Pi^{l+1}$ with respect to $\Pi^{l}$ with R.N. derivative proportional to $\Lambda(X)^{\psi_{l+1}-\psi_{l}}$. Instead of targeting $\Pi$ directly by $\mathbb{Q}$, one progressively steps through this sequence of artificial distributions, starting from $\mathbb{Q}$ to target the flattened $\Pi^{1}$ and ending up at targeting $\Pi$ with proposals drawn from $\Pi^{L-1}$.
@@
After $\Pi^{l+1}$ has been determined, particles are resampled based on the weights $\Lambda(X)^{\psi_{l+1}-\psi_{l}}$. Subsequently, to increase particle diversity, each particle is moved by a few steps of an MCMC scheme that targets $\Pi^{l+1}$. This can be done by proposing independent samples from $\mathbb{Q}$ and accepting/rejecting the proposals in a Metropolis-Hastings step based on the likelihood in Equation (35). Alternatively, proposals can be localised based on the preconditioned Crank-Nicolson scheme, see Section 4 for details. The complete tempering and move procedure is summarised in Algorithm 2. It is to be carried out at each observation time $t_{i}$ in place of Steps $2.$ and $3.$ in Algorithm 1.
Input: Samples $X^{(j)}$ of $\mathbb{Q}=\Pi_{i-1}\otimes\mathcal{L}^{g_{i}}$ with weights $\Lambda(X^{(j)})$, hyperparameters $\alpha$ and NN
@@ 3. Move particles:
For ease of exposition, the MCMC move step in Algorithm 2 is presented using independent proposals drawn from $X^{\prime}\sim\mathbb{Q}$. These proposals can be localised in Wiener space by using the preconditioned Crank-Nicolson scheme as is further explained in Section 4. For practical purposes, this implies having to store at the observation time $t_{i}$ for each particle $X^{(j)}$ the sample of the Wiener process $W$ that corresponds to simulating the sample path $X^{(j)}_{[t_{i-1},t_{i}]}$ following the dynamics in (32). Note that this is still carried out in an online setting since the stored Wiener paths relate to the evolution on $t_{i-1},t_{i}]$ only and may hence be discarded after each complete filtering step.
@@ 4. Guided smoothing via Metropolis-Hastings sampling
To localise the proposals drawn from $\mathcal{L}^{g}(X)$ we introduce random-walk-type proposals in the Wiener space by adapting a preconditioned Crank-Nicolson (pCN) scheme (see [50], [7], [21]) as follows.
@@
Input: Parameters A,F,QA,~F,~Q and $x_{0}$, observations $y$, iterations $N,N_{0}$, step size $\beta$
@@ 2. Save current state $X$
Since $W$ is a cylindrical Wiener process on $H$, it cannot be treated directly as an $H$-valued process. However, $W$ can be defined as a Wiener process taking values in a larger Hilbert space $H^{\prime}$ such that $H\hookrightarrow H^{\prime}$ is embedded in a Hilbert-Schmidt way. Note that such an embedding always exists, see [31], Section 4.4 for details. The solution map $\Gamma$ in (36) is then to be understood as a measurable mapping $\Gamma:H\times C([0,T];H^{\prime})\to C([0,T];H)$.
@@
Input: Parameters A,F,QA,~F,~Q and $\rho$, observations $y$, iterations $N,N_{0}$, step sizes $\beta,\beta_{0}$
@@ 4. Save current state $(V,X_{0},\theta)$
Remark $4.$
If AA and $Q$ are independent of the unknown parameter $\theta$, one only needs to solve the backwards ODEs $(U_{t})_{t\in[0,T]}$ and $(V_{t})_{t\in[0,T]}$ in (28) and (29) once. This severely eases the computational burden of Algorithm 5 for the highly relevant case that the parameter $\theta$ of interest parametrises only the nonlinearity $F=F_{\theta}$.
@@ 5.1. Stochastic Amari equation
with nonlinearity $F:L^{2}(D)\to L^{2}(D)$ defined by
@@
with respect to the Fourier basis functions $(e_{l})_{l=1}^{\infty}$ on $L^{2}(D)$. The operator $Q$ is then the Hilbert-Schmidt integral operator defined by the Matérn kernel on $D$, see for example [8], Theorem 5.
@@ 5.2. Filtering
Experiments are carried out on data generated by forward simulating the solution $X$ to Equation (42) on the spatiotemporal domain $0,T]\times D=[0,20]\times[-10\pi,10\pi]$ with step sizes $\Delta_{t}=0.02$ and $\Delta_{x}=20\pi/2^{8}$.
@@ Methods
    The guided particle filter as proposed in [44]. This method differs to ours in the proposal distribution $\mathbb{Q}$ of Eq. (31). The proposals chosen in [44] are defined by solving the guided process (32) with guiding term
@@
    The unscented Kalman filter (UKF) introduced in [68]. The UKF has been proposed for the filtering problem of an Amari model in [61], where it has been shown to perform well in the setting of a discretisation of Equation (42).
@@
Hence, once equilibrium is reached, the nonlinear part of the dynamical system vanishes, rendering a temporal discretisation more forgiving. As can be observed in Figure 3(d), with increasing time lag between observations, the performance of the UKF worsens drastically to the point of diverging in the case of $n=5$.
A comparison of the two particle filter methods in Figure 3(a) shows that GPF-II with proposals based on the guiding function in (48) slightly outperforms our proposed GPF-I in the dense observation setting. This may be explained yet again by the models convergence to an equilibrium state. Since the guiding function defined in (48) is derived based on the choice $\mathop{}\!\mathrm{d}Z_{t}=\sqrt{Q}\mathop{}\!\mathrm{d}W_{t}$ for the auxiliary process $Z$, once equilibrium is reached and $X$ behaves following (49), such proposals are almost locally optimal. Hence, GPF-II starts outperforming GPF-I around the observation times just before $t=8.0$ when an equilibrium state is ‘fully’ reached the first time. On the other hand, when observations are sparse, the quality of the proposals for the GPF-I at states away from equilibrium lead to a better performance in the cases of $n=10$ and $n=5$. Interestingly, as can be seen in Figure 3(b), the performance of GPF-I in this experiment increases as observations become more sparse. This coincides with findings in [44] for GPF-II in the context of a different dynamical system. In our experiment, this increase in performance is drastic enough for GPF-I in the $n=5$ case to be the best performing method over all observation schemes.
@@
For the second filtering experiment, the comparison between relative filtering errors of the three methods can be found in Figure 5. In this parametrisation with $\delta=0.5$, the system does not converge to an equilibrium; instead, it exhibits traveling wave behaviour. Expectedly, the proposed GPF-I outperforms GPF-II over all three observation schemes. We also see the UKF performing worse than the particle filter methods, even in the dense observation scheme. Moreover, the large time steps between observations plus the nonlinear dynamics of the traveling waves states result in significantly more divergent behaviour in the case of $n=5$ observations. On the other hand, GPF-II performs only slightly worse as observation numbers decrease, whereas GPF-I remains robust to the sparsity of observations.
@@ 5.3. Smoothing and parameter estimation
We apply Algorithm 5 to the given dataset with N=15  [^0] N=15\,000 MCMC iterations and pCN step size $\beta=0.1$. As priors we choose
@@
The parameters $\{\eta,\zeta,A,\delta\}$ are updated individually in a Gibbs sampling manner. Proposals for each parameter are chosen via an adaptive scaling Metropolis scheme (see e.g. [4]) as follows. At iteration $j$, given a step size $S^{\eta}_{j}$, say we propose $\eta_{j+1}$ via
@@ Results
Code availability:
The source code for the experiments in Section 5 is available at
@@
$\mathbb{E}[E^{h}_{t}\mid\mathcal{F}_{s}]$ $=\frac{1}{C^{h}}\mathbb{E}\left[\left(\prod_{j=1}^{i-1}k(X_{t_{j}},y_{j})\right)h(t,X_{t})\mid\mathcal{F}_{s}\right]$
@@
$\mathbb{E}[E^{h}_{t}\mid\mathcal{F}_{s}]$ $=\frac{1}{C^{h}}\mathbb{E}\left[\left(\prod_{j=1}^{i-1}k(X_{t_{j}},y_{j})\right)h(t,X_{t})\mid\mathcal{F}_{s}\right]$
@@
Let first $t\in[0,t_{1}]$ and denote by $K$ the infinitesimal space-time generator of the Markov process $X$. Following (50), it holds that $Kh=0$. From this and the Fréchet differentiability of $h$ in $x$ it follows with Lemma 3.3 in [53] that
@@
This, jointly with $E^{h}_{t}=h(t,X_{t})/C^{h},~t\in[0,t_{1}],$ shows (52) on $0,t_{1}]$. Now, let $t\in(t_{i-1},t_{i}],i\geq 2,$ and assume (52) holds for all $0,t_{i-1}].$
Again, by the definition of $E^{h}$, Equation (50) and Lemma 3.3 in [53] we have that
@@
Step Three. The general claim follows by repeating the backwards recursion of the first two steps to obtain $g(t,x)$ on the complete interval $0,t_{n}].$
@@
Hence, denoting by $K$ the infinitesimal space-time generator of $X$, it follows from [47], Theorem 4.1 that $Kg(t,x)=\langle F(t,x),\mathop{}\!\mathrm{D}_{x}g(t,x)\rangle$. Lemma 3.3 of [53] therefore gives that
@@
Consequently, plugging in $E^{g}$ as in (22) and applying the integration by parts for semimartingales we have on $0,t_{1})$ that
@@
This shows that $\bar{E}^{g}_{t}:=E^{g}_{t}/E^{g}_{0}=\mathcal{E}(M^{g})_{t}$ and hence (54) for all $t\in[0,t_{1}]$. The extension onto the complete time interval $[0,T]$ follows by the same inductive argument as in the proof of Theorem 2.4. Lastly, to show that $\mathcal{E}(M^{g})$ is a proper martingale, it suffices to note that $x\mapsto G(t,x)$ is Lipschitz continuous in $x$, uniformly on $[0,T]$. The martingale property of $\mathcal{E}(M^{g})$ then follows from [53], Lemma C.1. ∎
@@
This shows (56) and uniqueness of the mild solution follows from uniqueness of the infinite-dimensional forward Ricatti equation, see for example [10], Theorem 3.6 as well as [6], Chapter 2.2 and references within.
@@ B.1. Proofs of Section 5
Given Ax=xAx=-x, it holds that $(S_{t})_{t}$ is the semigroup given by $S_{\Delta_{i}}=\exp(-\Delta_{i})$. Hence, following the definition of $Q_{t}$ given in (15), we have
@@
Appendix C
The idea of guiding bears similarities to a data assimilation technique known as nudging, where a dynamical system described by a partial differential equation is adjusted by superimposing a data-dependent term. This “nudging” term is designed to “nudge” trajectories of the system towards observations. The underlying idea dates back to the 1960s and 1970s and is also known as Newtonian relaxation ([33]) and Luenberger observer in control theory for linear systems ([45]). The proposed nudging term is rather ad-hoc and consequently does not satisfy any minimum error criterion. Recently, [18] (see also [17]) derived nudging terms in a more principled way. In particular, the two proposed nudging terms proposed in [18] –PNDRQF (Physical Nudging Data-Retrieval Quadratic Form) and PNDRQ (Physical Nudging Data-Retrieval Quadrature)– are special instances of guiding terms inherited from a Gaussian system. In [63], a particle filter for SPDEs is introduced in which the nudging term is optimised to increase the effective sample size of the ensemble of particles.
The nudging approach differs from ours in several ways. Firstly, nudging is designed to tackle the filtering problem, and nudging terms are thereby restricted to one future observation. In contrast, the guiding terms we propose can be generalised to take all future observations into account, thereby enabling methodology to efficiently address the smoothing problem. Moreover, contrary to our approach, no Radon-Nikodym derivative for the law of the nudged process with respect to the true conditioned process is derived, precluding exact inference. Lastly, current approaches derive the nudging terms after discretising infinite-dimensional SDEs into a high-dimensional SDE. Guided processes, however, are derived in the fully infinite-dimensional Hilbert space setting, hence ensuring methodology that remains valid under mesh refinement.